# Towards Agentic Intelligence for Materials Science

### v1

@@ 3.7 Pipeline-Centric Perspective
Figure 7: Moving beyond proxy benchmarks in AI for materials science. The diagram highlights the limitations of current reactive models, which are often optimized for isolated tasks (proxies) like property prediction rather than real-world discovery. It illustrates the necessity of embedding these fine-tuned components into a broader agentic loop, where they serve as adaptive tools driven by end-to-end experimental outcomes rather than static dataset scores.
@@
After covering Q2, we discuss the interaction between the materials science fine-tuned LLMs and the agentic LLMs as shown in Fig. 1. On the one hand, existing pre-trained LLMs lack plasticity 98, 332, focusing on how these two components jointly implement decision-making in our pipeline-centric framework Moreover, compared to the general pre-training data, MatSci domain-specific data, as discussed in Secs. 3.1, 3.2, and 3.3, is orders of magnitude smaller even if we apply techniques like data augmentation as discussed in Sec. 3.5.1. These observations suggest that most behaviors of fine-tuned models are largely inherited from the pre-training stage, making it difficult to reliably elicit the desired behaviors using only limited MatSci-specific fine-tuning.
@@
Stepping back across the previous subsections, it becomes evident that most current AI4MatSci formulations are defined around narrowly scoped, mostly supervised tasks such as property prediction, information extraction, structure generation, and local optimization, each tied to bespoke datasets that are limited in both quantity and curation quality compared to the web-scale corpora used for general LLM pre-training as discussed in Sec. 2, and are typically optimized in isolation from end-to-end discovery outcomes. Even when augmented by multi-fidelity modeling, synthetic data, or automated extraction pipelines, these datasets remain sparse and noisy, making them at best partial proxies for the rich, open-ended discovery problem that materials science ultimately cares about, leaving a substantial gap between benchmark success and real experimental impact. A straightforward response would be to scale up the quality, diversity and quantity of supervised MatSci tasks and benchmarks. However, from an end-to-end, pipeline-centric perspective that explicitly targets the discovery of novel and useful materials under finite resources, over-optimizing models for performance on such intermediate tasks risks locking effort into local improvements on fragile surrogates (e.g., scores on a particular property dataset) while neglecting whether these gains actually translate into faster, safer identification and validation of new functional materials. In this view, the MatSci-tuned tasks surveyed in this section should be regarded as adjustable operators within a larger discovery pipeline, not as endpoints; their design, data, and evaluation criteria ought to be continuously re-examined through the lens of the ultimate discovery reward, so that limited high-quality MatSci data is allocated to those parts of the pipeline that most directly advance real experimental outcomes rather than merely inflating intermediate scores. Consequently, we argue that effort should gradually shift from reactive intermediate MatSci benchmarks, as discussed in this section, toward more agentic systems that explicitly close the loop with simulation and experiment, which will be developed in detail next.
@@
On the other hand, increasing plasticity of the pretraining or adaptation stage carries the risk of eroding fundamental safety and security properties. Materials-science-specific tasks cannot reasonably be expected to cover the full spectrum of normative constraints encoded in general-purpose pretraining corpora, so aggressive adaptation must be balanced against the need to preserve these foundational principles. From the pipeline-centric standpoint, the challenge is therefore to couple relatively scarce, noisy MatSci supervision with general-purpose pretraining in a way that steers the overall system toward the end-to-end objective of autonomous, safe materials discovery, without sacrificing the broad alignment and robustness inherited from the foundation model.
@@ 4 Agentic Systems for Materials Science
In this part, we try to address the guiding question Q3 in Sec. 1: What capacities have been developed, and how can we bridge the gap between existing AI for materials science and a fully end-to-end autonomous material discovery pipeline?
@@ 4.1 Contemporary Agentic Systems in Materials Discovery
Over the past ten years, AI’s role in materials science has changed from being a passive tool for analysis to an active participant in discovery. Early research mainly focused on prediction with graph neural networks (e.g., CGCNN 393, ALIGNN 79) and pre-trained models (e.g., CHGNet 94). They can learn the mappings between structures and properties from static datasets. These approaches showed that complex atomic relationships in materials could be encoded into high-dimensional latent representations. However, these AI models still operated passively. They could only predict but not interact. Therefore, agentic materials discovery establishes a new paradigm: AI is redefined as an autonomous scientific partner that capable of independently generating hypotheses, planning experiments, and continuously learning through feedback loops227, 46.
An agentic system in materials research operates according to the principle of closed-loop discovery. Formally, let a material’s state be represented by a descriptor vector $x\in\mathbb{R}^{n}$, with measurable properties $y=f(x)$ obtained through experiment or simulation. Traditional supervised learning seeks to approximate $f$. In contrast, an agentic system defines a control policy $\pi_{\theta}(a|s)$ that selects actions $a$ (e.g., synthesis parameters, composition proposals) based on system state $s$ (e.g., prior experimental results, model uncertainty) to maximize a long-term objective $R=\mathbb{E}[\sum_{t}\gamma^{t}r_{t}]$. In materials settings, this reward typically balances experimental performance, resource cost, and safety considerations rather than only predictive accuracy. Reinforcement learning (RL) thus becomes a natural formalism for autonomous experimentation 35. However, in practice, the state and action spaces are highly non-convex, expensive to explore, and embedded with scientific constraints such as thermodynamic stability or synthesis feasibility. Consequently, the success of agentic systems depends on the integration of uncertainty quantification, active learning, and surrogate modeling to guide exploration efficiently247, 15. This integration distinguishes fully agentic pipelines 28 from simpler Bayesian optimization loops by coupling rich state representations.
@@
Self-driving laboratories (SDLs) play an important role in autonomous loops. An SDL combines robotic hardware for synthesis and characterization with AI models that predict outcomes, evaluate uncertainty, and iteratively plan the next experiment 343. In these systems, the learning algorithm closes the “materials loop”: (1) Design and selection: propose an experiment or select programs in the database, (2) Execution: perform synthesis autonomously, (3) Evaluation: measure outcomes, and (4) Evolution: retrain the model on updated data. This cycle implements concrete credit assignment from measured outcomes back to the models that propose and prioritize experiments. Studies in autonomous thin-film optimization 258 and catalyst discovery 354 have demonstrated that such loops can simplify experimental search spaces and preserve physical interpretability at the same time. AI’s job in autonomous loops shifts from passive reasoning to agentic cognition with the ability to autonomously generate and evaluate hypotheses.
Recently, hierarchical and multi-agent architectures have been developed to extend this autonomy recently. For example, MatExpert 96 utilized multiply stages to mimic the human materials design expert’s workflow. In multi-agent system design, there is no longer just one single decision-making entity. The discovery process is decomposed into several specialized agents, such as Predictor, Planner, and Verifier. Those agents can communicate with each other via standardized protocols 123. The Predictor utilizes neural networks (e.g., CHGNet 94, ALIGNN 79) to evaluate candidate materials, while the Planner uses Bayesian optimization or RL to choose the promising directions in the action space. The Verifiers are embodied by robotic automation and are capable of conducting experiments and sending the results back to the shared memory. These architectures enable modular expansion and enhance explainability. Each agent follows a well-defined scientific rubric and simulates the division of human labor within research teams.
@@
Agentic systems must also balance exploration and exploitation. Pure exploitation may converge to local optima because it continuously refines known successful materials, while unbounded exploration wastes too many resources on random new paths. Techniques such as upper confidence bound (UCB) sampling and Thompson sampling in Bayesian optimization frameworks 327 provide theoretical guarantees for efficient exploration. Meanwhile, recent RL approaches integrate epistemic uncertainty from ensemble models 202 to quantify the expected information gain of each new experiment. In materials discovery, this translates into a quantifiable trade-off between novel compositions and reliable performance. Such uncertainty-aware strategies embody the concept of “safe exploration” in the physical sciences40.
Another important development is the emergence of simulation and experiment in hybrid simulation-experiment loops173. High-fidelity simulations, including density functional theory (DFT) or molecular dynamics (MD) can act as virtual laboratories, generating synthetic data to pre-train models or provide experimental priors. Frameworks like DeepMD 168 and CHGNet 94 provide fast computation that makes it capable of running millions of virtual experiments before a single synthesis is attempted. Agentic systems that combine simulation with experiment bridge the gap between computation and reality. Virtual agents can co-evolve with their physical counterparts, which is closely related to the concept of digital twins for materials discovery 177.
The future trajectory of agentic systems will likely involve scientific multi-agent ecosystems336, in which autonomous entities collaborate, compete, and refine hypotheses and ideas under shared goals. By combining large language models (LLMs) for reasoning and documentation with reinforcement learning planners and simulation engines, these ecosystems could perform discovery , hypothesis explanation, literature grounding, and experimental justification in natural language363, 14, 357. Such “AI laboratories” will require governance mechanisms320 to ensure safety and reproducibility, but they are also capable of compressing the scientific feedback cycle from months to hours.
@@ 4.2 Scientist AI: Beyond Data Fitting to Scientific Reasoning
The new phase in AI for materials science goes from making accurate predictions to using AI to do scientific reasoning on its own. This means systematically acquiring, testing, and changing beliefs about the physical world353. Early deep learning frameworks such as CGCNN 393 and ALIGNN 79, discussed in Sec. 4.1, are capable of learning intricate structure–property mappings. However, they cannot articulate hypotheses, mechanisms, or explanatory frameworks353.
In contrast, a scientist AI system 372 aims to operationalize scientific reasoning. It independently formulates hypotheses, orchestrates virtual or tangible experiments, and updates its internal models in response to the resultant evidence 227.Instead of just guessing what attributes would happen with certain inputs, this kind of system keeps and improves a list of possible explanations for how structure, processing, and environment affect behavior353. It also looks for data that will help test and improve those explanations353. Previous research on robot scientists 189 and self-driving laboratories 355 illustrates aspects of this concept, wherein automated systems repetitively suggest experiments, gather data, and revise internal models accordingly. This sub-section delineates emerging methodologies—active learning, autonomous experimentation, symbolic inference, and literature-grounded reasoning—as approximating various elements of a reasoning-centric workflow. Nonetheless, these elements have yet to be amalgamated into a cohesive scientific methodology: hypotheses are infrequently articulated, causal or mechanical reasoning is limited, and the integration of priors across modalities remains fragile.
@@ 4.2.1 Hypothesis Generation and Search Over Scientific Spaces
One of Scientist AI’s main skills is using priors, evidence, and uncertainty to recommend pathways in the materials space353. Predictive models can inform practitioners what the attributes of a target will be based on certain inputs. However, making hypotheses demands working in organized scientific domains where composition, processing parameters, mechanisms, and physical restrictions all matter353. Recent advancements in machine learning and autonomous testing exhibit significant potential for proposal generation; yet, they still lag far behind human scientific thinking367.
@@
Contemporary Bayesian optimization (BO) and active learning frameworks have exhibited robustness in suggesting useful experiments characterized by quantified uncertainty. These methods usually use Gaussian process surrogates or similar probabilistic models 105. The goal is to find a balance between exploration and exploitation in high-dimensional design spaces 289. In materials science, BO has facilitated the swift optimization of alloys, catalysts, and thin-film systems367. It helps reduce the number of trials by several orders of magnitude and allowing for closed-loop experimental workflows 200. These systems partially implement the hypothesis-proposal phase of scientific reasoning by identifying promising areas inside the design space amongst existing uncertainty41.
@@
Generative models offer an alternative approach to hypothesis formulation. Variational autoencoders, generative adversarial networks, and diffusion-based models have been utilized in molecular and crystalline design 23, yielding candidate structures tailored to certain target properties or design goals 256. These methods increase the effective hypothesis space, allowing for the investigation of chemically valid yet previously unrecognized compositions and microstructures. Recent research illustrate inverse-design workflows wherein generative models provide candidates that fulfill various objectives or constraints, including stability and performance thresholds, succeeded by evaluation using high-fidelity simulations or experiments 23.
@@
LLMs add another way to come up with hypotheses. LLMs and science-specific variations, like SciBERT, can read a lot of scientific texts and make sense of them37. They can also find domain-specific entities and relationships and provide mechanistic narratives or qualitative design rationales89. Recent agentic systems that link LLMs to chemistry and materials tools demonstrate that these models can suggest reaction pathways, enumerate viable synthetic routes, and advocate for experimental parameter sweeps by utilizing patterns found in the literature and outputs from external tools47. These systems are increasingly including explicit priors and limitations when used with retrieval and organized knowledge bases89. This makes it easier to come up with hypotheses that are based on current materials knowledge and makes it easier to integrate with that knowledge 297.
@@
Even with these improvements, today’s systems still do not use hypothesis generation in a scientifically sound way. First, hypotheses are seldom articulated openly; the majority of approaches provide candidate points or structures without delineating the corresponding mechanical assumptions, validity regimes, or expected failure mechanisms367. Second, generative proposals are frequently assessed solely via surrogate property predictors, exhibiting minimal linkage to falsifiable assertions or established theories, hence limiting their function in systematic hypothesis testing and refinement367. Third, constructing long-horizon hypotheses—like multi-step reasoning from mechanism to candidate design to experimental protocol—continues to be fragile and heavily reliant on limited training distributions or manually generated prompts18. Lastly, even though LLM-based agents seem to be able to combine literature and coordinate tools, their factual reliability, calibration, and grounding in domain knowledge are still problems that need to be addressed. This is shown in broader evaluations of foundation models 49 and domain-specific LLMs for chemistry 297.
@@
All of these changes show that AI systems may suggest candidate materials, designs, and mechanistic theories based on data and with an awareness of uncertainty353. Nonetheless, they have not yet produced organized, falsifiable hypotheses or engaged in systematic reasoning on alternative explanatory theories18. To improve this skill, generative modeling has to be more closely linked to physical limitations, clear hypothesis representations, and systematic uncertainty quantification within a larger agentic loop for autonomous materials discovery353.

### v2

@@ 3.7 Pipeline-Centric Perspective
After covering Q2, we discuss the interaction between the materials science fine-tuned LLMs and the agentic LLMs as shown in Fig. 1. On the one hand, existing pre-trained LLMs lack plasticity 99, 334, focusing on how these two components jointly implement decision-making in our pipeline-centric framework Moreover, compared to the general pre-training data, MatSci domain-specific data, as discussed in Secs. 3.1, 3.2, and 3.3, is orders of magnitude smaller even if we apply techniques like data augmentation as discussed in Sec. 3.5.1. These observations suggest that most behaviors of fine-tuned models are largely inherited from the pre-training stage, making it difficult to reliably elicit the desired behaviors using only limited MatSci-specific fine-tuning.
@@
Stepping back across the previous subsections, it becomes evident that most current AI4Mat-Sci formulations are defined around narrowly scoped, mostly supervised tasks such as property prediction, information extraction, structure generation, and local optimization, each tied to bespoke datasets that are limited in both quantity and curation quality compared to the web-scale corpora used for general LLM pre-training as discussed in Sec. 2, and are typically optimized in isolation from end-to-end discovery outcomes. Even when augmented by multi-fidelity modeling, synthetic data, or automated extraction pipelines, these datasets remain sparse and noisy, making them at best partial proxies for the rich, open-ended discovery problem that materials science ultimately cares about, leaving a substantial gap between benchmark success and real experimental impact. A straightforward response would be to scale up the quality, diversity and quantity of supervised MatSci tasks and benchmarks. However, from an end-to-end, pipeline-centric perspective that explicitly targets the discovery of novel and useful materials under finite resources, over-optimizing models for performance on such intermediate tasks risks locking effort into local improvements on fragile surrogates (e.g., scores on a particular property dataset) while neglecting whether these gains actually translate into faster, safer identification and validation of new functional materials. In this view, the MatSci-tuned tasks surveyed in this section should be regarded as adjustable operators within a larger discovery pipeline, not as endpoints; their design, data, and evaluation criteria ought to be continuously re-examined through the lens of the ultimate discovery reward, so that limited high-quality MatSci data is allocated to those parts of the pipeline that most directly advance real experimental outcomes rather than merely inflating intermediate scores. Consequently, we argue that effort should gradually shift from reactive intermediate MatSci benchmarks, as discussed in this section, toward more agentic systems that explicitly close the loop with simulation and experiment, which will be developed in detail next.
@@
On the other hand, increasing plasticity of the pre-training or adaptation stage carries the risk of eroding fundamental safety and security properties. Materials-science-specific tasks cannot reasonably be expected to cover the full spectrum of normative constraints encoded in general-purpose pre-training corpora, so aggressive adaptation must be balanced against the need to preserve these foundational principles. From the pipeline-centric standpoint, the challenge is therefore to couple relatively scarce, noisy MatSci supervision with general-purpose pre-training in a way that steers the overall system toward the end-to-end objective of autonomous, safe materials discovery, without sacrificing the broad alignment and robustness inherited from the foundation models.
@@ 4 Agentic Systems for Materials Science
In this part, we try to address the guiding question Q3 in Sec. 1: What capacities have been developed, and how can we bridge the gap between existing AI for materials science and a fully end-to-end autonomous materials discovery pipeline?
@@ 4.1 Contemporary Agentic Systems in Materials Discovery
Over the past ten years, AI’s role in materials science has changed from being a passive tool for analysis to an active participant in discovery. Early research mainly focused on prediction with graph neural networks (e.g., CGCNN 396, ALIGNN 80) and pre-trained models (e.g., CHGNet 95). They can learn the mappings between structures and properties from static datasets. These approaches showed that complex atomic relationships in materials could be encoded into high-dimensional latent representations. However, these AI models still operated passively. They could only predict but not interact. Therefore, agentic materials discovery establishes a new paradigm: AI is redefined as an autonomous scientific partner that capable of independently generating hypotheses, planning experiments, and continuously learning through feedback loops229, 47.
An agentic system in materials research operates according to the principle of closed-loop discovery. Formally, let a material’s state be represented by a descriptor vector $x\in\mathbb{R}^{n}$, with measurable properties $y=f(x)$ obtained through experiment or simulation. Traditional supervised learning seeks to approximate $f$. In contrast, an agentic system defines a control policy $\pi_{\theta}(a|s)$ that selects actions $a$ (e.g., synthesis parameters, composition proposals) based on system state $s$ (e.g., prior experimental results, model uncertainty) to maximize a long-term objective $R=\mathbb{E}[\sum_{t}\gamma^{t}r_{t}]$. In materials settings, this reward typically balances experimental performance, resource cost, and safety considerations rather than only predictive accuracy. RL thus becomes a natural formalism for autonomous experimentation 36. However, in practice, the state and action spaces are highly non-convex, expensive to explore, and embedded with scientific constraints such as thermodynamic stability or synthesis feasibility. Consequently, the success of agentic systems depends on the integration of uncertainty quantification, active learning, and surrogate modeling to guide exploration efficiently249, 15. This integration distinguishes fully agentic pipelines 29 from simpler Bayesian optimization loops by coupling rich state representations.
@@
Self-driving laboratories (SDLs) play an important role in autonomous loops. An SDL combines robotic hardware for synthesis and characterization with AI models that predict outcomes, evaluate uncertainty, and iteratively plan the next experiment 345. In these systems, the learning algorithm closes the “materials loop”: (1) Design and selection: propose an experiment or select programs in the database, (2) Execution: perform synthesis autonomously, (3) Evaluation: measure outcomes, and (4) Evolution: retrain the model on updated data. This cycle implements concrete credit assignment from measured outcomes back to the models that propose and prioritize experiments. Studies in autonomous thin-film optimization 260 and catalyst discovery 357 have demonstrated that such loops can simplify experimental search spaces and preserve physical interpretability at the same time. AI’s job in autonomous loops shifts from passive reasoning to agentic cognition with the ability to autonomously generate and evaluate hypotheses.
Recently, hierarchical and multi-agent architectures have been developed to extend this autonomy recently. For example, MatExpert 97 utilized multiply stages to mimic the human materials design expert’s workflow. In multi-agent system design, there is no longer just one single decision-making entity. The discovery process is decomposed into several specialized agents, such as Predictor, Planner, and Verifier. Those agents can communicate with each other via standardized protocols 124. The Predictor utilizes neural networks (e.g., CHGNet 95, ALIGNN 80) to evaluate candidate materials, while the Planner uses Bayesian optimization or RL to choose the promising directions in the action space. The Verifiers are embodied by robotic automation and are capable of conducting experiments and sending the results back to the shared memory. These architectures enable modular expansion and enhance explainability. Each agent follows a well-defined scientific rubric and simulates the division of human labor within research teams.
@@
Agentic systems must also balance exploration and exploitation. Pure exploitation may converge to local optima because it continuously refines known successful materials, while unbounded exploration wastes too many resources on random new paths. Techniques such as upper confidence bound (UCB) sampling and Thompson sampling in Bayesian optimization frameworks 329 provide theoretical guarantees for efficient exploration. Meanwhile, recent RL approaches integrate epistemic uncertainty from ensemble models 203 to quantify the expected information gain of each new experiment. In materials discovery, this translates into a quantifiable trade-off between novel compositions and reliable performance. Such uncertainty-aware strategies embody the concept of “safe exploration” in the physical sciences41.
Another important development is the emergence of simulation and experiment in hybrid simulation-experiment loops174. High-fidelity simulations, including density functional theory (DFT) or molecular dynamics (MD) can act as virtual laboratories, generating synthetic data to pre-train models or provide experimental priors. Frameworks like DeepMD 169 and CHGNet 95 provide fast computation that makes it capable of running millions of virtual experiments before a single synthesis is attempted. Agentic systems that combine simulation with experiment bridge the gap between computation and reality. Virtual agents can co-evolve with their physical counterparts, which is closely related to the concept of digital twins for materials discovery 178.
The future trajectory of agentic systems will likely involve scientific multi-agent ecosystems338, in which autonomous entities collaborate, compete, and refine hypotheses and ideas under shared goals. By combining LLMs for reasoning and documentation with RL planners and simulation engines, these ecosystems could perform discovery, hypothesis explanation, literature grounding, and experimental justification in natural language366, 14, 360. Such “AI laboratories” will require governance mechanisms322 to ensure safety and reproducibility, but they are also capable of compressing the scientific feedback cycle from months to hours.
@@ 4.2 Scientist AI: Beyond Data Fitting to Scientific Reasoning
The new phase in AI for materials science goes from making accurate predictions to using AI to do scientific reasoning on its own. This means systematically acquiring, testing, and changing beliefs about the physical world355. Early deep learning frameworks such as CGCNN 396 and ALIGNN 80, discussed in Sec. 4.1, are capable of learning intricate structure–property mappings. However, they cannot articulate hypotheses, mechanisms, or explanatory frameworks355.
In contrast, a scientist AI system 375 aims to operationalize scientific reasoning. It independently formulates hypotheses, orchestrates virtual or tangible experiments, and updates its internal models in response to the resultant evidence 229.Instead of just guessing what attributes would happen with certain inputs, this kind of system keeps and improves a list of possible explanations for how structure, processing, and environment affect behavior355. It also looks for data that will help test and improve those explanations355. Previous research on robot scientists 190 and self-driving laboratories 358 illustrates aspects of this concept, wherein automated systems repetitively suggest experiments, gather data, and revise internal models accordingly. This sub-section delineates emerging methodologies—active learning, autonomous experimentation, symbolic inference, and literature-grounded reasoning—as approximating various elements of a reasoning-centric workflow. Nonetheless, these elements have yet to be amalgamated into a cohesive scientific methodology: hypotheses are infrequently articulated, causal or mechanical reasoning is limited, and the integration of priors across modalities remains fragile.
@@ 4.2.1 Hypothesis Generation and Search Over Scientific Spaces
One of Scientist AI’s main skills is using priors, evidence, and uncertainty to recommend pathways in the materials space355. Predictive models can inform practitioners what the attributes of a target will be based on certain inputs. However, making hypotheses demands working in organized scientific domains where composition, processing parameters, mechanisms, and physical restrictions all matter355. Recent advancements in machine learning and autonomous testing exhibit significant potential for proposal generation; yet, they still lag far behind human scientific thinking370.
@@
Contemporary Bayesian optimization (BO) and active learning frameworks have exhibited robustness in suggesting useful experiments characterized by quantified uncertainty. These methods usually use Gaussian process surrogates or similar probabilistic models 106. The goal is to find a balance between exploration and exploitation in high-dimensional design spaces 291. In materials science, BO has facilitated the swift optimization of alloys, catalysts, and thin-film systems370. It helps reduce the number of trials by several orders of magnitude and allowing for closed-loop experimental workflows 201. These systems partially implement the hypothesis-proposal phase of scientific reasoning by identifying promising areas inside the design space amongst existing uncertainty42.
@@
Generative models offer an alternative approach to hypothesis formulation. Variational autoencoders, generative adversarial networks, and diffusion-based models have been utilized in molecular and crystalline design 24, yielding candidate structures tailored to certain target properties or design goals 258. These methods increase the effective hypothesis space, allowing for the investigation of chemically valid yet previously unrecognized compositions and microstructures. Recent research illustrate inverse-design workflows wherein generative models provide candidates that fulfill various objectives or constraints, including stability and performance thresholds, succeeded by evaluation using high-fidelity simulations or experiments 24.
@@
LLMs add another way to come up with hypotheses. LLMs and science-specific variations, like SciBERT, can read a lot of scientific texts and make sense of them38. They can also find domain-specific entities and relationships and provide mechanistic narratives or qualitative design rationales90. Recent agentic systems that link LLMs to chemistry and materials tools demonstrate that these models can suggest reaction pathways, enumerate viable synthetic routes, and advocate for experimental parameter sweeps by utilizing patterns found in the literature and outputs from external tools48. These systems are increasingly including explicit priors and limitations when used with retrieval and organized knowledge bases90. This makes it easier to come up with hypotheses that are based on current materials knowledge and makes it easier to integrate with that knowledge 299.
@@
Even with these improvements, today’s systems still do not use hypothesis generation in a scientifically sound way. First, hypotheses are seldom articulated openly; the majority of approaches provide candidate points or structures without delineating the corresponding mechanical assumptions, validity regimes, or expected failure mechanisms370. Second, generative proposals are frequently assessed solely via surrogate property predictors, exhibiting minimal linkage to falsifiable assertions or established theories, hence limiting their function in systematic hypothesis testing and refinement370. Third, constructing long-horizon hypotheses—like multi-step reasoning from mechanism to candidate design to experimental protocol—continues to be fragile and heavily reliant on limited training distributions or manually generated prompts18. Lastly, even though LLM-based agents seem to be able to combine literature and coordinate tools, their factual reliability, calibration, and grounding in domain knowledge are still problems that need to be addressed. This is shown in broader evaluations of foundation models 50 and domain-specific LLMs for chemistry 299.
@@
All of these changes show that AI systems may suggest candidate materials, designs, and mechanistic theories based on data and with an awareness of uncertainty355. Nonetheless, they have not yet produced organized, falsifiable hypotheses or engaged in systematic reasoning on alternative explanatory theories18. To improve this skill, generative modeling has to be more closely linked to physical limitations, clear hypothesis representations, and systematic uncertainty quantification within a larger agentic loop for autonomous materials discovery355.